OpenStudy (anonymous):

s=n/2[2a+(n-1)d] make a as the subject of the formula medals will be given

OpenStudy (anonymous):

a=−1/2dn+1/2d+n/4s

OpenStudy (anonymous):

how can we give medals

OpenStudy (anonymous):

by clicking best response

OpenStudy (perl):

sheelo, can you post your last question again, i think i can solve it now

OpenStudy (perl):

the question with C = x^2 + y^2 -

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

C=x^2+y^2-z^2\[\div 2xy\]

OpenStudy (perl):

solve for z ?

OpenStudy (anonymous):

yep

OpenStudy (perl):